OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x. x f(x) = 9x + 7 g(x) = 5x -2 -11 -1 -2 0 1 1 5 2 Complete the chart and determine the solution of the equation f(x) = g(x)? x = -1 x = 0 x = 2 x = 25

OpenStudy (the_fizicx99):

You can either graph them and put the value at which they intersect OR plug in the x values for all the x's and see which gives you a true statement For example: 6 = 6 is a true statement, therefore a solution <-- example

OpenStudy (anonymous):

how do I plug in whered the -11 and -2 come from where do I figure out the other terms to fill in

OpenStudy (the_fizicx99):

Huh?

OpenStudy (anonymous):

how am I supposed to fill in the chart

OpenStudy (anonymous):

for g(x), where x is -2, put -2 in the equation for g(x)

OpenStudy (anonymous):

g(x) = 5x. Put -2 in the equation. 5(-2) = -10

OpenStudy (anonymous):

so g(-2) = -10

OpenStudy (anonymous):

oh I just notice the 5x has an exponent of x so g(x) = 5x ^x so it would be = 5?

OpenStudy (anonymous):

for x=-2

OpenStudy (anonymous):

I'm back

OpenStudy (anonymous):

5x^x (-2)^(-2) = -(1/4). 5*(-1/4) = (-5/4)

OpenStudy (anonymous):

internet trouble

OpenStudy (anonymous):

I cant write a fraction into the table it wouldn't make sense

OpenStudy (anonymous):

Why not? (-5/4) = -1.25 if you want to use decimals

OpenStudy (anonymous):

can you do the rest yourself?

OpenStudy (anonymous):

wait how does -2^-2 equals -1/4 wouldn't it equal 4

OpenStudy (the_fizicx99):

An exponent can't be negative it would be 1/-2^2 = 1/4 = 0.25

OpenStudy (anonymous):

sorry, it's positive (1/4) for -2^-2

OpenStudy (anonymous):

1/(-2)^2

OpenStudy (anonymous):

so the answer is (5/4) or 1.25

OpenStudy (anonymous):

I just used a calculator and -2^2 = 4

OpenStudy (anonymous):

did you put brackets in between -2, like this: (-2)^(-2). if you do -2^-2 on a calculator, it will give you the wrong number.

OpenStudy (the_fizicx99):

All I saw is that you were asking if -2^-2 is 4, i said no because an exponent can't be negative, therefore to make it positive you make it 1/-2^2 = 1/4 = 0.25 IF it was -2^2 it is 4 because -2 * -2 = 4 as the two negatives cancel out and make a positive

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

ok so if x= -1 I would do 1/-1^1=1 over -1

OpenStudy (the_fizicx99):

No, ONLY IF THE EXPONENT IS NEGATIVE

OpenStudy (anonymous):

but if x is the exponent and you need to plug in -1 wouldn't the exponent be negative

OpenStudy (the_fizicx99):

I don't know what you're trying to solve, but just know that any exponent that is negative must have a 1/a^b to make it positive. From what I see they aren't exponents, 5x and 9x isn't being rasied to any power.. that would be x^5 or x^9

OpenStudy (anonymous):

im trying to fill out the table

OpenStudy (the_fizicx99):

Screen shot it and make a new question, its too messy. I'll help you

OpenStudy (anonymous):

ok